OpenStudy (faman39):

Which of the following is true? 302 F < 150C 302 F > 150 C 302 F = 150 C 302 C < 150 F

OpenStudy (faman39):

@LukeBlueFive i was looking for other but its been a few hours, no one came to help, thats y, i m forced to call you , becuz you helped more than i need and i do not wanted bother more today, but now, i am helpless

OpenStudy (faman39):

Can you help me with this?

OpenStudy (anonymous):

\[150\times1.8+32=302\] The third one looks like the right one.

OpenStudy (anonymous):

Just to be clear, convert F° into C° °C = ( 5/9 ) ( °F - 32 ) = ( 5/9 ) ( 302 - 32 ) = ...